Please help my daughter with the following problem:
The second of two numbers is 4 more than the first. The sum is 56. find the numbers.

Answer by pwac(253)

make x the first number
so second number is x+4.This gives following sum:
x+x+4=56 add left side to give:
2x+4=56 subtract 4 from both sides:
2x=52 divide both sides by 2
x=26 this is your first number
x+4=30 which is second number
26+30=56
hope this helps
Pete